# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v1

@@ Introduction
Evaluation Methodology
Robustness. Robustness reflects the classifier’s ability to maintain its performance beyond training examples. We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [52, 50]. Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) and fools the model into making incorrect predictions. The p-norm of the perturbation vector is limited by $\epsilon$ to ensure the perturbation is imperceptible.
We craft adversarial examples using three gradient-based adversarial algorithms, FGSM [40], PGD [50], and DeepFool [53], with details introduced in Section A.1 in the Appendix. These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image. For conventional classifiers, we directly optimize the perturbations using three types of adversarial algorithms. However, for VLM-based classifiers (VLM+RoBERTa), the standard optimization approach is not applicable because we lose continuous gradient flow after VLMs generate word tokens. To address this challenge, we approximate the gradients by instructing the VLMs to generate a target output ($y_{tar}$), which is contrary to the image label. For example, if the VLM initially classifies a safe image with the output “This image is safe,” we compute the gradients when the target output is “unsafe.” Different from conventional classifiers, where we maximize the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$). We craft adversarial examples with FGSM and PGD for open-source VLM-based classifiers (LLaVA and InstructBLIP). Note, we exclude DeepFool as it requires the gradients of logits, which are not directly applicable in this context.
Evaluation Metric for Robustness. We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers. RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images. Here, we generate perturbed images only for those that are correctly predicted by classifiers during the effectiveness evaluation. Furthermore, we calculate the Mean Square Error (MSE) between original images and adversarial examples to reveal the magnitude of perturbations required to mislead classifiers into making incorrect predictions.
Evaluation Results
@@
Classifier-Level. The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V. It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception. GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738). Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784. However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories. In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.
@@
Different Unsafe Categories. To understand which unsafe categories are more effectively detected, we count the number of classifiers capable of identifying each category and record the average F1-Score in Figure 2. We find that the Sexual and Shocking categories have higher average F1-Scores, close to 0.8, compared to other unsafe categories. They are also covered by more classifiers, with 5-7 classifiers capable of identifying such content. In contrast, the remaining unsafe categories, particularly Spam, Hate, Harassment, and Self-Harm, have lower average F1-Scores (below 0.5). This discrepancy highlights the imbalanced moderation across different types of unsafe content.
@@
Real-World Versus AI-Generated. Real-world images from LAION-5B and AI-generated ones from Lexica have different distributions (see 11(a) in the Appendix). To compare the performance of tested classifiers on two groups of images, we calculate the average F1-Score of each classifier on real-world and AI-generated images in Figure 3. We find that SD_Filter, NSFW_Detector, NudeNet, and GPT-4V have better performance in identifying real-world unsafe images compared to AI-generated ones. For example, the F1-Score of SD_Filter decreases from 0.833 on real-world images to 0.727 on AI-generated images. Meanwhile, MultiHeaded, LLaVA, and InstructBLIP have higher F1-Scores on AI-generated images than real-world ones. The different behaviors of classifiers on two sets of images are multifaceted. A key factor is the distribution shift between training and testing images, including nuanced differences in semantics, image styles, etc. After examining the training images of the classifiers (if available), we find that classifiers like NudeNet and NSFW_Detector are trained on real-world NSFW images [18], as indicated by [17] and [16]. Meanwhile, the training data for SD_Filter and GPT-4V has not been disclosed. For classifiers that are more effective on AI-generated images, i.e., MultiHeaded, we find that it is trained on images generated by text-to-image models.
@@
Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it. We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images. The first group consists of SD_Filter, NFSW_Detector, and NudeNet. These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category. However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4. The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images. To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica. To characterize images from these two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics. Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information. We then apply KMeans clustering with $K$ ranging from 2 to 15 and employ the Elbow method [^13] https://www.scikit-yb.org/en/latest/api/cluster/elbow.html. to determine the optimal number of clusters. This method shows that the optimal number of clusters is 5-8 depending on different classifiers. Finally, we create five clusters and retrieve the nearest image to each cluster centroid. We display these central images, representing their respective clusters in Figure 4 and Figure 13 in the Appendix, each figure corresponding to one group of classifiers.
@@
Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout. First, images in an artistic representation, [^14] It is an open question whether unsafe content in artistic representation remains unsafe. In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form. despite showing explicitly unsafe content, can often escape detection of classifiers. For example, among false negatives shown in 4(a), AI-generated images within clusters 1, 2, 4, and 5 showcase nudity in an artistic style, and yet they are misclassified as safe. In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset. Similarly, among the misclassified images from the Violence category (see 13(a) in the Appendix), AI-generated images within clusters 1, 2, and 4 are also misclassified, which portray artistic representations of violence, such as fighting scenes. Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers. These images are composites made of multiple smaller images, commonly four or nine, arranged in a grid layout. Examples include cluster 2, 3, 5 in 4(a), cluster 5 in 4(b), and cluster 2, 4 in 13(a). Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the performance degradation, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
@@
To test this hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions. To this end, we randomly select 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories. These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only. For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [60]. Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images. By adjusting the strength parameter, we control the degree of the artistic style applied to the original image. For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6. Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid. Figure 5 shows image variations with an example from the Violence category. In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations. To guarantee that the variations in the artistic representation retain the same content as the original images, we conducted a manual annotation to filter out any variations where the content has been altered. We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16. We present the percentage of correct predictions in Figure 6. The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ performances generally decrease. This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
@@
Table 5: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations. We report the MSE (average square error) and RA (robust accuracy).
GN FGSM PGD DeepFool Average
Model MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$)
Q16 <1e-3 0.972 <1e-3 0.402 <1e-3 0.004 <1e-3 0.160 <1e-3 0.385
MultiHeaded <1e-3 0.988 <1e-3 0.596 <1e-3 0.012 <1e-3 0.218 <1e-3 0.454
SD_Filter <1e-3 1.000 <1e-3 0.552 <1e-3 0.000 <1e-3 0.876 <1e-3 0.607
NSFW_Detector <1e-3 0.976 <1e-3 0.496 <1e-3 0.012 <1e-3 0.372 <1e-3 0.464
NudeNet <1e-3 1.000 <1e-3 0.086 <1e-3 0.004 <1e-3 0.084 <1e-3 0.293
LLaVA* <1e-3 0.996 1.365 0.710 0.849 0.180 - - 0.738 0.629
InstructBLIP* <1e-3 0.996 1.417 0.624 0.881 0.204 - - 0.766 0.608
@@
Take-Aways. We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories. Our findings reveal several insights. First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content. However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets. Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content. Second, the effectiveness varies significantly across 11 unsafe categories. Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement. For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V. Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images. The underlying reason is the distribution shift between these two types of images. Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and the composition of unsafe images in a grid layout. These distinct characteristics may pose challenges to classifiers trained only on real-world images, leading to the observed performance degradation.
@@
Setup. Generating adversarial examples for the entire UnsafeBench dataset can be time-consuming, because it requires performing an independent optimization process for each image. To address this issue, we randomly sample 500 images that each classifier correctly predicted to evaluate their robustness. We then generate perturbed images by adding Gaussian noise and creating optimized perturbations using three types of adversarial algorithms: FGSM, PGD, and DeepFool. We employ the $L_{\infty}$ norm to constrain perturbations and set $\epsilon$ to 0.01 for all experiments. We demonstrate the perturbed images and different types of perturbations in Figure 7. With these perturbed images, we measure the MSE and RA (metrics are introduced in Section 3.4) of seven open-source classifiers. Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires access to the model’s internal parameters, which is not available for GPT-4V.
Figure 7: Perturbed images with Gaussian perturbations and three types of adversarial perturbations. Adversarial examples generated with FGSM, PGD, and DeepFool cause Q16 to misclassify these unsafe images as safe.
@@
Results. Table 5 demonstrates the MSE and RA of seven open-source classifiers across four types of perturbations. VLM-based classifiers show the highest robustness. They achieve RAs between 0.629 and 0.608, higher than any conventional classifiers. This suggests that a significant percentage of perturbed images can still be correctly classified by VLM-based classifiers. Moreover, their MSEs range from 0.738 to 0.766, far exceeding the conventional classifiers, whose MSEs are less than $31e-$ 3. This indicates that it requires significant perturbations to deceive VLM-based classifiers. Meanwhile, Q16, MultiHeaded, SD_Filter, and NSFW_Detector present a moderate level of robustness, with RAs ranging from 0.385 to 0.607. However, NudeNet shows the lowest RA, i.e., 0.293, revealing that it is the most vulnerable classifier to adversarial perturbations. Even imperceptible perturbations with MSE less than $31e-$ 3 can deceive NudeNet in most cases involving adversarial examples.
@@
To explore different levels of robustness, we review the model architecture, training paradigm, and training dataset for each classifier (see Section 2). As the least robust classifier, NudeNet is an Xception-based classifier and is trained on 160K labeled images using fully supervised learning. Other conventional classifiers are built on CLIP, which is pre-trained on 400 million image-text pairs and then fine-tuned on their own labeled datasets using linear probing or prompt learning. They present higher robustness compared to training a small classifier using supervised learning, i.e., NudeNet. VLM-based classifiers present the highest robustness and also rely on large pre-trained models. For example, LLaVA utilizes CLIP as the image feature extractor and LLaMA [68] as the LLM to reason over the input image and generate output. To conclude, we find that under the attack of random perturbed and adversarial examples, relying on large pre-trained foundational models enhances robustness compared to training new, smaller classifiers from scratch. For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
@@
The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs. First, their effectiveness needs to be improved, particularly in several neglected categories such as Hate and Self-harm. Second, these classifiers face challenges in detecting AI-generated unsafe images due to distributions that differ from real-world images. With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content. Using this dataset, we take the first step towards building a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI. PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain. The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity. Similarly, PerspectiveVision evaluates the safety of a given image and further identifies the specific category if the image is deemed unsafe. Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content. The high-level overview is demonstrated in Figure 8.
@@
Table 6: F1-Score of PerspectiveVision models on the UnsafeBench test set.
Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
Linear Probing (CLIP) 0.776 0.736 0.767 0.586 0.924 0.894 0.877 0.841 0.907 0.903 0.899 0.859
Prompt Learning (CLIP) 0.437 0.557 0.679 0.337 0.890 0.758 0.765 0.596 0.829 0.635 0.654 0.687
LoRA Fine-Tuning (LLaVA) 0.686 0.755 0.760 0.743 0.913 0.861 0.889 0.780 0.878 0.909 0.855 0.844
@@
SMID MultiHeaded-D NudeNet-D Hate Symbols Self-Harm Violent Behaviors Overall
Linear Probing (CLIP) 0.365 0.654 0.954 0.598 0.801 0.740 0.723
Prompt Learning (CLIP) 0.623 0.500 0.984 0.849 0.971 0.931 0.810
LoRA Fine-Tuning (LLaVA) 0.557 0.675 0.986 0.593 0.948 0.908 0.794
GPT-4V 0.532 0.577 0.991 0.786 0.924 0.921 0.818

### v2

@@ Introduction
Conventional vs. VLM-Based Classifiers. The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V. It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception. GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738). Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784. However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories. In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.
@@
Imbalanced Effectiveness Across Different Categories. To understand which unsafe categories are more effectively detected, we count the number of classifiers capable of identifying each category and record the average F1-Score in Figure 3. We find that the Sexual and Shocking categories have higher average F1-Scores, close to 0.8, compared to other unsafe categories. They are also covered by more (5-7) classifiers. In contrast, the remaining unsafe categories, particularly Hate, Harassment, and Self-Harm, have lower average F1-Scores (below 0.5). This discrepancy highlights the imbalanced effectiveness across different types of unsafe content.
@@
Real-World vs. AI-Generated. Real-world images from LAION-5B and AI-generated ones from Lexica have different distributions. To compare the performance of tested classifiers on two groups of images, we calculate the average F1-Score of each classifier on real-world and AI-generated images in Figure 4. We find that SD_Filter, NSFW_Detector, NudeNet, and GPT-4V have better performance in identifying real-world unsafe images compared to AI-generated ones. For example, the F1-Score of SD_Filter decreases from 0.833 on real-world images to 0.727 on AI-generated images. Meanwhile, MultiHeaded, LLaVA, and InstructBLIP have higher F1-Scores on AI-generated images than real-world ones. The different behaviors of classifiers on two sets of images are multifaceted. A key factor is the distribution shift between training and testing images, including nuanced differences in semantics, image styles, etc. Recall the training images of the classifiers provided in Table 1, classifiers like NudeNet and NSFW_Detector are trained on real-world NSFW images [16], as indicated by [15] and [14]. Moreover, MultiHeaded is trained only on unsafe images by text-to-image models, thus demonstrating better performance on AI-generated images.
@@
Effectiveness Degradation of Classifiers Trained on Real-World Images Only. Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it. We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images. The first group consists of SD_Filter, NFSW_Detector, and NudeNet. These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category. However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4. The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images. To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica. To characterize images from two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics. Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information. We then apply KMeans clustering with $K=4$ for better visualization purposes. Finally, we create four clusters for each group of images and retrieve the nearest image to each cluster centroid. We display these central images, representing their respective clusters in Figure 5 and Figure 14 in the Appendix, each figure corresponding to one group of classifiers.
@@
Distinct Characteristics in AI-Generated Images. Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout. First, images in an artistic representation, [^5] It is an open question whether unsafe content in artistic representation remains unsafe. In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form. despite showing explicitly unsafe content, can often escape detection of classifiers. For example, among false negatives shown in 5(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe. In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset. Similarly, among the misclassified images from the Violence category (see 14(a) in the Appendix), AI-generated images within clusters 1, 2, and 3 are also misclassified, which portray artistic representations of violence, e.g., fighting scenes. Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers. These images are composites made of smaller images, commonly four or nine, arranged in a grid layout. Examples include cluster 1, 3 in 5(a), and cluster 3 in 14(a). Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
@@
Case Study on Artistic Representation and Grid Layout. To test our hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions. To this end, we randomly sample 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories. These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only. For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [57]. Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images. By adjusting the strength parameter, we control the degree of the artistic style applied to the original image. For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6. Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid. Figure 6 shows image variations with an example from the Violence category. In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations. To guarantee that the variations in the artistic representation retain the same content as the original images, we conduct a manual annotation to filter out any variations where the content has been altered. We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16. We present the percentage of correct predictions in Figure 7. The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ effectiveness generally decreases. This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
@@
Table 5: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations. We report the mean robust accuracy (RA) and the standard deviation based on results from three sampling times. All perturbations are bounded by a fixed $\epsilon=0.01$ and obtained within the same number of iterations (1 for GN/FGSM; 100 for PGD/DeepFool).
Dataset Model GN FGSM PGD DeepFool Average
LAION-5B Q16 1.000 $\pm$0.000 0.497 $\pm$0.003 0.002 $\pm$0.002 0.170 $\pm$0.011 0.417 $\pm$0.002
MultiHeaded 1.000 $\pm$0.000 0.707 $\pm$0.025 0.001 $\pm$0.002 0.321 $\pm$0.040 0.507 $\pm$0.016
SD_Filter 1.000 $\pm$0.000 0.603 $\pm$0.001 $0$$\pm$0.000 0.047 $\pm$0.011 0.413 $\pm$0.003
NSFW_Detector 0.999 $\pm$0.001 0.603 $\pm$0.005 0.001 $\pm$0.001 0.423 $\pm$0.020 0.507 $\pm$0.006
NudeNet 1.000 $\pm$0.000 0.105 $\pm$0.013 $0$$\pm$0.000 0.092 $\pm$0.009 0.299 $\pm$0.006
LLaVA* 0.945 $\pm$0.011 0.784 $\pm$0.020 0.712 $\pm$0.005 0.458 $\pm$0.010 0.725 $\pm$0.006
InstructBLIP* 0.995 $\pm$0.002 0.621 $\pm$0.003 0.472 $\pm$0.013 0.163 $\pm$0.006 0.563 $\pm$0.006
Lexica Q16 0.999 $\pm$0.001 0.313 $\pm$0.016 0.001 $\pm$0.001 0.115 $\pm$0.016 0.357 $\pm$0.008
MultiHeaded 0.999 $\pm$0.001 0.474 $\pm$0.008 0.001 $\pm$0.001 0.157 $\pm$0.004 0.408 $\pm$0.001
SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 $0$$\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 $0$$\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
NudeNet 1.000 $\pm$0.000 0.076 $\pm$0.000 0.003 $\pm$0.000 0.082 $\pm$0.000 0.290 $\pm$0.000
LLaVA* 0.959 $\pm$0.016 0.648 $\pm$0.009 0.587 $\pm$0.015 0.469 $\pm$0.023 0.666 $\pm$0.010
InstructBLIP* 0.999 $\pm$0.001 0.648 $\pm$0.017 0.565 $\pm$0.001 0.324 $\pm$0.015 0.634 $\pm$0.007
@@
Take-Aways. We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories. Our findings reveal several insights. First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content. However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets. Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content. Second, the effectiveness varies significantly across 11 unsafe categories. Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement. For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V. Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images. Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and unsafe images in a grid layout.
@@
Results. We randomly sample 500 images three times that are correctly classified by each classifier and create adversarial examples. Table 5 lists the RA of seven open-source classifiers for four types of perturbations. VLM-based classifiers show the highest robustness. They achieve RAs between 0.563-0.725 on LAION-5B images and 0.634-0.666 on Lexica images, higher than any conventional classifiers. Meanwhile, conventional classifiers are more vulnerable to adversarial attacks, with RAs of 0.299-0.507 on LAION-5B images and 0.290-0.444 on Lexica images. Among them, NudeNet shows the lowest RA (0.290-0.299), revealing that it is the most vulnerable classifier to adversarial attacks.
@@
Relying on Pre-Trained Foundational Models Enhances Robustness Than Training Smaller Classifiers From Scratch. We review the model architecture, training paradigm, and training dataset for each classifier (see Section 2). As the least robust classifier, NudeNet is an Xception-based classifier and is trained on 160K labeled images using fully supervised learning. Other conventional classifiers are built on CLIP, which is pre-trained on 400 million image-text pairs and then fine-tuned on their own labeled datasets using linear probing or prompt learning. They present higher robustness compared to training a small classifier using supervised learning, i.e., NudeNet. VLM-based classifiers present the highest robustness and also rely on large pre-trained models. For example, LLaVA utilizes CLIP as the image feature extractor and LLaMA [65] as the LLM to reason over the input image and generate output. For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
Figure 8: Confidence scores of classifiers’ predictions over different groups of images. The median value is denoted next to the dashed line. AI-generated images tend to have lower confidence scores than real-world images.
Classifiers Are More Susceptible to Adversarial Attacks using AI-Generated Images Than Real-World Images. We find that most classifiers, except for InstructBLIP, exhibit lower RAs on AI-generated images compared to their real-world counterparts. Their average RAs show a consistent decrease from the range of 0.299–0.725 to 0.290–0.666, with a maximum drop of 10%. To investigate the reason behind this, we calculate the maximum probability as the confidence score of four conventional classifiers when classifying different groups of images. As shown in Figure 8, successful adversarial examples (denoted in blue) generally have lower confidence scores. This indicates that these examples are closer to the decision boundary and can be more easily perturbed to cross over this boundary compared to non-adversarial examples. When comparing AI-generated and real-world images, AI-generated images tend to have lower confidence scores in the distribution. This potentially contributes to the higher robustness susceptibility for these classifiers.
Interestingly, we also find these classifiers only show evident RA decreases on adversarial perturbations (especially FGSM), and not with Gaussian perturbations. The difference in creating the two types of perturbations is that adversarial perturbations are designed to maximize the training loss, i.e., cross-entropy loss, whereas Gaussian perturbations are not designed to do so. We then analyze classifiers’ cross-entropy loss and their loss change due to the addition of adversarial perturbations, using FGSM as an example. In Figure 15 (in the Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones. This implies that classifiers tend to be more sensitive to the perturbations in AI-generated images, leading to a higher loss increase even with the same amount of perturbation. This also explains why, even if the median confidence scores and loss values are close between AI-generated images and real-world images, such as the NSFW_Detector, the AI image group still makes the classifier more vulnerable to adversarial attacks, because they are more prone to crossing the decision boundary.
@@
The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs. First, there is a lack of an open-source image moderation tool that effectively identifies the wide spectrum of unsafe images, particularly in several neglected categories such as hateful content and self-harm. Second, these classifiers show both degraded effectiveness and robustness in detecting AI-generated unsafe images due to distributions that differ from real-world images. With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content. Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI. PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain. The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity. Similarly, PerspectiveVision is designed to evaluate the safety of a given image and further identify the specific category if the image is deemed unsafe. Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content. The high-level overview is demonstrated in Figure 9.
@@
In-Distribution Out-of-Distribution (OOD)
Classifier UnsafeBench-Test SMID MultiHeaded-D NudeNet-D Self-Harm Violent Behaviors OOD Average
Linear Probing (CLIP) 0.859 0.365 0.654 0.954 0.801 0.740 0.752
Prompt Learning (CLIP) 0.687 0.623 0.500 0.984 0.971 0.931 0.802
LoRA Fine-Tuning (LLaVA) 0.844 0.557 0.675 0.986 0.948 0.908 0.836
GPT-4V 0.717 0.532 0.577 0.991 0.924 0.921 0.830